a proton has an electrical potential energy of 2x10^-16 J relative to a nearby negatively charged plat. IF the proton is released from rest and accelerates towards the negative plate, how fast will it be going when it reaches the plate?
 
 

Grade:12th pass

2 Answers

Eshan
5 years ago
Dear student,

When the proton reaches the negatively charged plate, it loses all its electrical potential energy into its kinetic energy(from Law of Conservation of Energy).

Hence\dfrac{1}{2}mv^2=2\times 10^{-16}

\implies v=5\times 10^5m/s
Khimraj
5 years ago
The velocity will be given by
KE = poential energy
½mv2 = 2*10-16
So v = 5*105m/s
Hope it clears..............................